OpenStudy (anonymous):

Solve the system: 2x + 8y = 1 x = 2y

Parth (parthkohli):

Hello there, as x = 2y, we can write x as 2y in the first equation. \( \color{Black}{\Rightarrow 2(2y) + 8y = 1 }\) \( \color{Black}{\Rightarrow 4y + 8y = 1}\)

Parth (parthkohli):

Can you solve for \(y\)?

OpenStudy (anonymous):

\(x\) is whatever you get when you replace \(y\) in either equation by what you get for \(y\)

Parth (parthkohli):

Find \(x\) after solving for \(y\). Since x is two times y, just multiply 2 to y for getting x.

OpenStudy (anonymous):

umm 12y=1?

Parth (parthkohli):

Yep

Parth (parthkohli):

Now divide both sides by 12, and you get y = 1/12

OpenStudy (anonymous):

y=1?

Parth (parthkohli):

What do you get for x? You multiply two to y, so 1/12 * 2

Parth (parthkohli):

12y = 1 1y = 1/12 Find x by multiplying 2 to y

OpenStudy (anonymous):

Plug 2y into the first equation for x. 2(2y) + 8y = 1 4y + 8y = 1 12y = 1 y = 1/12 Now to find x, substitute 1/12 into one of the equations for y. x = 2(1/12) x = 1/6

OpenStudy (anonymous):

y = 1/12 x = 1/6 :)